OpenStudy (anonymous):

Write an equation in slope-intercept form for a line perpendicular to the graph of the equation that passes through the x-intercept of that line y=-1/2x-4 The answer to this is y=2x+16 and I don't know how the book got there, please explain.

OpenStudy (anonymous):

Okay. We know that a perpendicular slope to an original function is the normal line of the original function's slope. Opposite reciprocal. If the slope is 2x, the perpendicular line's slope is -1/2x. If the slope is 2/5 x, the perpendicular line's slope is -5/2x. If the slope, in this case, is -1/2x, then the perpendicular line's slope is 2/1x. So, if the equation is y=-1/2x -4, then the perpendicular line will be y=(2/1x )+some number. Mathematically, what did it take to get from -1/2 to 2/1? Multiplied by negative four (other than the opposite reciprocal trick.) So your constant will be multiplied by -4. -4 times -4 = 16. So your equation is y=(2/1)x + 16.

OpenStudy (anonymous):

Thank you! @hink